ACTIVITY #4



Scoring system: 1 point for correct answer and 2 points for the solution



A. Use the Table of Standard Normal distribution to determine corresponding probabilities of the following:



1. P(Z < -1.45)



2. P(0.5<Z < 2.33)



3. P(Z > 1.78)



B. X is a normally distributed random variable with a mean of 60 and a standard deviation of 8. Find the probabilities indicated by using the table.



1. P(X < 52)



2. P(48 < X < 64)



3. P(X > 57)



C. Solve the following problems by supplying the needed information.



The average number of calories in a 1.8-ounce chocolate bar is 230. Suppose that the distribution of calories is approximately normal with a standard deviation of 10.



1. Find the probability that a randomly selected chocolate bar will have less than 200 calories.



2. Find the probability that a randomly selected chocolate bar will have greater than 195 calories. 3. What is the percentage that a chocolate bar is randomly selected between 200 calories and 250 calories?

1
Expert's answer
2022-04-06T07:22:15-0400

Using the standard Z table

A.

  1. P(Z<-1.45)=0.0735
  2. P(0.5<Z<2.33)=0.9901-0.6915=0.2986
  3. P(Z>1.78)=0.0375

B.

Z="\frac{x-\mu }{\sigma}"

sample mean=60

standard deviation =8

1.P(x<52)

z="\frac {52-60}{8}"=-1

P(z"\le"-1)= 0.1587

Thus P(x<52) is 0.1587.

2.P(48<x<64)=

P(z1<z<z2)

z1="\frac{48-60}{8}" =-1.5

z2="\frac{64-60}{8}" =1.5

P(z1<z<z2)=P(z"\le" 1.5)-P(z"\le" -1.5)

=0.9332-0.0668

=0.8864

3.P(x>57)=1-P(x"\le"57)

z="\frac{57-60}{8
}"

z=-0.375

P(z>57)=1-0.3538

P(z>57)=0.6462

C.

mean =230

s=10

Z="\frac{x-\mu }{\sigma}"

1.

P(x<200)=P("z=\frac{200-230}{10})"

P(z <-3)=0.0014

 The probability that a randomly selected chocolate bar will have less than 200 calories is 0.0014.


2.

P(x >195)=1-P(x<195)

P(x>195)=1-P("z\le\frac{195-230}{10})"

P(x>195)=1-P("z\le
-3.5)"

=1-0.0002

=0.9998

 The probability that a randomly selected chocolate bar will more than 195 is 0.9998.

3.

P(200<x<250)=P(z1<z<z2)

z1="\frac{200-230}{10} =-3"

z2="\frac{250-230}{10}" =2

P(200<x<250)=P(z2=2)-P(z1=-3)

=0.9773-0.0014

=0.9760

The percentage that a chocolate bar is randomly selected between 200 calories and 250 calories is 97.60.